OpenStudy (anonymous):

What is the solution to the equation 4 2x ≈ 3 ? answers: x = -0.631 x = -0.396 x = 0.396 x = 0.631

OpenStudy (anonymous):



OpenStudy (anonymous):

4 2x ≈ 3 ? ^ unclear

OpenStudy (anonymous):

what do you mean?

OpenStudy (anonymous):

0.396

OpenStudy (anonymous):

what you actually mean is 4 ^(2x) ≈ 3

OpenStudy (campbell_st):

\[4^{2x} = 3\] take the base 4 log of each side and use change of base \[2x = \frac{\ln(3)}{\ln(4)}\] divide both sides by 2 \[x=\frac{1}{2}\times \frac{\ln(3)}{\ln(4)}\] just evalaute for the answer

OpenStudy (anonymous):

ohh ok thank you :]